OpenStudy (anonymous):

The diagnal of a rectangle measures 45 cm. The Width of the rectangle measures27cm. Find the Length of the rectangle. THe length of the rectangle is?

OpenStudy (anonymous):

L = sqrt ( 45^2 - 27^2) = sqrt ( 1296) = 36

OpenStudy (anonymous):

Thanks. Did you use the Pythagorean theory (formula)

OpenStudy (anonymous):

Yes use pythagorean theorem with the diagonal being the hypotenuse

OpenStudy (anonymous):

ok, thanks